# RatBodyFormer: Rodent Body Surface from Keypoints

### v1

@@ Abstract
Rat behavior modeling goes to the heart of many scientific studies, yet the textureless body surface evades automatic analysis as it literally has no keypoints that detectors can find. The movement of the body surface, however, is a rich source of information for deciphering the rat behavior. We introduce two key contributions to automatically recover densely 3D sampled rat body surface points, passively. The first is RatDome, a novel multi-camera system for rat behavior capture, and a large-scale dataset captured with it that consists of pairs of 3D keypoints and 3D body surface points. The second is RatBodyFormer, a novel network to transform detected keypoints to 3D body surface points. RatBodyFormer is agnostic to the exact locations of the 3D body surface points in the training data and is trained with masked-learning. We experimentally validate our framework with a number of real-world experiments. Our results collectively serve as a novel foundation for automated rat behavior analysis and will likely have far-reaching implications for biomedical and neuroscientific research.
Figure 1: We introduce a novel multiview camera system (RatDome) to capture multiview videos of rats (top), and a novel Transfomer-based network (RatBodyFormer) that learns to predict dense body surface points of a rat from its detectable keypoints from large-scale data (RatDome Dataset) captured with it (bottom). Past methods predict only sparse 2D keypoints such as their faces and legs (orange points). Our RatBodyFormer turns them into the dense 3D body surface (colorful points) offering a much richer window into the complex rat behavior.
@@ 1 Introduction
Rodent behavior analysis underpins the scientific process of many areas in biomedical and neuroscientific research. The behavioral outcomes of rodents play a key role in validating hypotheses that lead to scientific discoveries. Careful observations of rodent behaviors, however, incur cost in manpower and time, causing a bottleneck in the scientific process. Human observation is also prone to errors. Subtle behavioral differences may be missed or misjudged [9, 31]. If we can automate this experimental validation process while ensuring its accuracy or even extend it beyond the human level, we may dramatically expedite the scientific discovery process [15, 33, 12, 8]. For instance, it would enable concurrent testings of many hypotheses.
@@
We focus on rats, the most commonly used rodent in scientific experiments. A growing number of recent works, indeed, target computational rat behavior analysis or provide the basis for it. Belongie et al. [2] pioneered rat tracking by modeling the whole body as a spatio-temporal blob. Chaumont et al. [7] refined the method by dividing each blob into three parts linked with physical constraints. More recently, Mathis et al. [1] introduced a pre-trained deep network that can be fine-tuned to extract 2D keypoints of interest. This provides a general framework for skeleton estimation of animals, very much in the spirit of human pose estimation [6].
@@
These past methods, however, only provide us with the means to understand rat behavior through the movements of sparse keypoints. With DeepLabCut [1], trackable points of a rat would typically lie on the face, hands, feet, and tail. Although the movements of these keypoints may suffice for a range of applications, for instance, to recognize when a rat is feeding itself, they only give a sparse sampling of the entire body which tells us much less than what the detailed movements of other body parts may convey.
@@
The largest missed opportunity lies in the body surface, the fur-coated elastic body that shows subtle twitches, twists, curl-up, stretching, and even hair standing which all eloquently speak to the rich inner conditions of the rat and colorizes their behavior with context. Missing the body surface is inevitable with a keypoint-based method, even with large-scale pre-training, as the rat body lacks any texture whatsoever for such keypoint detectors to latch on to. It is just a uniform white or dark-brown surface. It is also highly non-rigid and deformable, much more so than the human body. At the same time, even though the rats are the same at the DNA level, their body surfaces can be slightly different in size and shape across different ages in months.
@@
How then can we model the 3D body surface of rats? Our goal is to realize full 3D rat reconstruction including the body surface without interfering with its natural behavior. We achieve this by learning to recover the 3D body surface from the handful of keypoints that can be automatically detected. Our key idea is to learn and leverage the coordination between the well-defined keypoints and the non-rigid movements of the body surface.
@@
We make two key technical contributions for this. The first is a novel means to collect training data. We need a sizable amount of training data that associates representative keypoints including facial features and appendages with densely sampled points of the body surface. The rat body surface is textureless and deformable, to the point that it is near impossible even for a human to annotate. We overcome this by temporary attaching trackable points (colorful beads) and passively capturing them in a novel multi-camera system which we refer to as the RatDome. We show that we can automatically recover the keypoints together with the 3D body surface points at each time instance with multiview geometry. We call this first-of-its-kind dataset, the RatDome dataset. The second is a novel network to transform detected 3D keypoints to 3D body surface points. We realize this with a transformer-based model that is agnostic to the exact locations of the 3D body surface points in the training data and introduce masked-learning to train it. We refer to this model as RatBodyFormer.
@@
We experimentally validate our framework with a number of real-world experiments using rats for neuroscientific studies. We first show that our RatBodyFormer can estimate the body surface accurately regardless of their poses or shapes. Our model can estimate their body surface with an average L2 error of around 6mm, approximately the same diameter as the beads attached to the body surface. We also show that rat body movements can be predicted by simply forecasting the keypoints from their past movements. This may help better segregate the movements of multiple rodents and serve as a useful tool to understand their interactions. We believe our RatDome dataset and RatBodyFormer offer foundational tools for advancing the automation of rodent behavior analysis and together open a new avenue of research towards computer vision for science.
@@ 2 Related Work
Markerless, non-invasive methods have also been proposed [2, 35, 7, 5]. Belongie et al. [2] introduced a detection and tracking method for a mouse using spatio-temporal volumes of monocular video. Matsumoto et al. [35] tracked major body parts (head, neck, trunk, and hip) with 3D skeleton estimation with depth sensing.
More recently, deep learning based keypoint detection has become popular [39, 3, 1, 48, 10, 14, 46, 19]. DeepLabCut [1] detects manually specified keypoints through transfer learning of 2D human pose estimation from monocular images [40]. DANNCE [10] and FreiPose3D [48] similarly apply human pose estimation and integrate multiview images to estimate 3D keypoints. LiftPose3D [14] achieves monocular 3D animal pose estimation by leveraging a deep neural network for lifting 2D human poses to 3D [34].
All these methods only detect sparse keypoints such as facial points and paws and cannot be applied to the feature-less body surface [11]. Our method complements these methods by establishing a framework to extrapolate dense points of the textureless body surface from feature-rich keypoints completely noninvasively without any markers.
@@
One common approach for 3D human reconstruction is to learn a statistical model from pre-acquired 3D scans to better condition the solution space [30, 49, 51, 28]. SMPL [30] is a statistical 3D human linear blended shape model parameterized by the person’s shape and pose. The pose parameters are the joint rotations and the shape parameters are the PCA coefficients of a large collection of aligned body scans. SMAL [49] extends this to quadrupled animals. Most 3D reconstruction methods regress these parameters directly from the image with a neural network [23, 47] or optimize their parameters so that projected keypoints align with their imaged ones [4, 21, 25, 20]. Joints of rodents are, however, not well-defined as they are deeply veiled by their fluffy body. Instead, we leverage reliably detectable 3D keypoints to recover the body surface.
Recent works have also introduced image-based animal appearance modeling [42, 29, 50, 45]. AnimalAvatars [42] regresses continuous surface embedding [38] from an image to associate each pixel with a point on the target 3D shape in its canonical pose. This CSE estimation, however, requires large-scale dense annotation between 2D image pixels and a 3D canonical surface ala DensePose [16], which is costly and unreliable for textureless body surfaces. In contrast, we leverage beads and multiview geometry to learn the sparse keypoint to dense body surface mapping, which ensures positional consistency and robust mapping.
@@ 3 Method
We learn to estimate the rat body surface from its detected keypoints. First, we collect a dataset that associates densely sampled body surface points with detectable keypoints with a newly developed multiview camera system which we refer to as the RatDome (Sec. 3.1). Second, we derive a transformer-based network that takes the keypoint 3D coordinates as input and outputs sampled body surface point 3D coordinates. We refer to this network as RatBodyFormer (Sec. 3.2). The model is trained with the data collected with RatDome (RatDome Dataset).
@@ 3.1 RatDome Dataset
Figure 2: A color-beaded rat. We attach color (red, black, orange, blue) beads (left) and paint on the rat body surface (right).
Figure 3: RatDome is a novel multiview camera studio for freely moving rats. It is shaped as a 15-faced gyroelongated pentagonal pyramid. With 15 cameras and their multiview geometry, we collect, annotate, and reconstruct paired sets of 3D keypoints and 3D body surface points of rats of different ages in weeks.
Regardless of the pre-training, a network can only learn to detect points that exhibit sufficient visual features. As a result, even after fine-tuning, an off-the-shelf deep keypoint detector [1] can only reliably detect points on the face and appendages of the rat. Our goal is to learn to extrapolate the body surface from these keypoints. For this, we will need a sufficiently large-scale dataset of paired sets of 3D keypoints and 3D body surface points. This is challenging as the rat body is completely textureless.
We overcome this challenge by simply endowing the rat with body texture. As shown in Fig. 2, we attach color beads to the body as well as paint markers in the areas where beads cannot be attached. We refer to these two types of body markers, beads and paints, simply as markers. It is important to note that these markers are only used for training data capture and the rats are completely in their natural form when their behavior is observed in actual experiments. These markers are also small enough that they do not alter the rat’s behavior as far as we could tell.
We build a novel multi-camera system to passively observe the color-beaded rat and reconstruct the 3D coordinates of individual markers. As shown in Fig. 3, this RatDome is a rat-scale multiview studio. Its shape is a gyroelongated pentagonal pyramid with 15 faces. Each face is a equilateral triangle with sides of 400mm, and accommodates up to three cameras or microphones on it while also serving as a green background of the system. For our data capture, we mount one camera on each face totaling 15 views. RatDome follows the modular design of CMU Panoptic studio [22] and can adapt to new capture devices by replacing each faces.
By capturing freely moving rats of different ages (7, 9, 11 week-old) in RatDome, we collect multiview videos and the paired sets of 3D keypoints and 3D markers. We recorded 2 sessions for the 7 week-old, 1 session for the 9 week-old, and 2 sessions for the 11 week-old. Each session was approximately 10 to 15 minutes. We manually annotate the marker IDs for the 7 and 11 week-old rats for about 1100 frames each. We semi-automatically annotate (see Sec. 3.2) the 7, 9, and 11 week-old rats’ marker IDs for about 8000, 1900, and 10000 frames, respectively. The 7, 9, 11 week-old rats have 58, 37, and 66 surface markers, respectively. We refer to this first-of-its-kind large-scale rat body surface dataset as the RatDome Dataset.
@@ 3.2 RatBodyFormer
RatBodyFormer takes $K$ 3D keypoint coordinates as input and outputs $N$ 3D surface point coordinates. The keypoints are chosen as points locatable in images, and we use the following 10 points: nose, right and left eyes, ears, front paws, back paws, and the base of tail. We use these in lieu of body joints, e.g., shoulders, since the body joints are ill-defined from the appearance of a rat unlike humans or horses [30, 51, 49].
@@
To represent different body shapes of different rats, we introduce point-wise scaling and translation parameters for individual rat, $\mathbf{C}_{\mathrm{P}}=\{c_{p_{i}}\}_{1}^{K},\mathbf{T}_{\mathrm{P}}=\{t_{p_{i}}\}_{1}^{K}$ and $\mathbf{C}_{\mathrm{B}}=\{c_{b_{j}}\}_{1}^{N},\mathbf{T}_{\mathrm{B}}=\{t_{b_{j}}\}_{1}^{N}$ for keypoints and body surface points, respectively. These scaling and translation parameters are optimized in the training process. At inference time for a rat whose scaling and translation parameters are unknown, we test-time optimize the parameters while keeping RatBodyFormer frozen.
The keypoints and the body surface points are first rotated around the Z-axis, i.e., the direction of gravity, such that the rat face is oriented to a pre-determined direction. Their 3D coordinates are then normalized to $[-1:1]$: $\mathbf{P}=\{p_{i}\}_{1}^{K}$ and $\mathbf{B}=\{b_{j}\}_{1}^{N}$. The same normalization is also applied to $\tilde{S}$.
As shown in Fig. 4, the decoder outputs are the estimated normalized displacements of the body surface points from their coordinates in the reference pose. The decoder queries are normalized displacements of the body surface points in an initial guess $\hat{\mathbf{B}}$ from their coordinates in the reference pose. The initial guess is calculated analytically. We use ARAP [43] and deform $\tilde{S}$ with the constraint that $\tilde{\mathbf{P}}$ is aligned with $\mathbf{P}/\mathbf{C}_{\mathrm{P}}$. As a result, we obtain the ARAP-deformed $\tilde{\mathbf{B}}$ as $\hat{\mathbf{B}}=\{\hat{b}_{j}\}_{1}^{N}$. For each keypoint and body surface point, we use the displacement vector normalized by the identity-dependent parameters $p_{i}/c_{p_{i}}-(\tilde{p}_{i}+t_{p_{i}})$ and $\hat{b}_{j}/c_{b_{j}}-(\tilde{b}_{j}+t_{b_{j}})$ as the inputs to the model, respectively.
The encoder and the decoder first embed each of these displacements in higher dimension $d$ together with positional encoding. The positional encoding for keypoints and body surface points is defined with a set of sinusoidal functions of the projection of their positions $\tilde{p}_{i}$ and $\tilde{b}_{j}$ on $\tilde{S}$ by the Laplace-Beltrami eigenfunctions defined on the canonical body surface $\tilde{S}$ [38, 27]. The output of the RatBodyFormer decoder are $N$ normalized displacement vectors $\{\delta_{j}\}_{1}^{N}$. Each $\delta_{j}$ reconstructs the body surface coordinates as $c_{j}(\delta_{j}+\tilde{b}_{j}+t_{b_{j}})$. We set $d=128$ in the experiment.
Figure 4: RatBodyFormer is an encoder-decoder Transformer model that takes the normalized displacements of detected 3D keypoints and outputs the normalized discplacements of densely sampled 3D body surface points. The displacements are with respect to the reference pose.
@@
The primary loss function for training RatBodyFormer is the supervision provided by the annotations in the RatDome Dataset. We directly compare the estimated 3D body surface point coordinates and their ground-truth coordinates with L2 norm as $\mathcal{L}_{\mathrm{3D}}$. This L2 norm loss $\mathcal{L}_{\mathrm{3D}}$, however, can be defined only for a subset of the $N$ body surface points. Since they are defined as the union of all annotated body surface points on different rat surfaces, a single input keypoint set from a rat has a ground-truth annotation only for the body surface points of the same rat. Even for the body surface points from the same rat, some of them can be missing in the captured data due to occlusion.
In addition to the L2 norm loss $\mathcal{L}_{\mathrm{3D}}$, we employ a silhouette loss $\mathcal{L}_{\mathrm{s}}$ which counts the number of predicted body surface points whose projections fall outside of the 2D rat region in a view [17]
$\mathcal{L}_{\mathrm{s}}=\sum_{n=1}^{N}M_{c}\left(P_{c}(\bm{x}_{n})\right)\,,$ (1)
where $P_{c}$ is a projection function onto the view $c$ and $M_{c}$ is a mask image of view $c$. $M_{c}(\bm{v})$ is a pixel value of $M_{c}$ at the image coordinate $\bm{v}$.
@@
RatDome Dataset contains many frames capturing color-beaded rats but without manual annotations (Sec. 3.1). We derive a semi-supervised learning method to make full use of this data. The method leverages a small number of frames fully-annotated with ground-truth to infer the labels for the remaining frames. First, we compute the 3D coordinates of the markers. This process is automated, similar to that of the Panoptic Studio Dataset [22]. We detect the 2D markers from each image with an object detector [13] trained with manual annotations. We then leverage multiview triangulation to disambiguate markers of the same color on the body surface.
Once the 3D marker coordinates are computed for each frame, we assign their marker IDs so that they are consistent with the manually annotated ones by using RatBodyFormer itself. Suppose we have trained RatBodyFormer using only the manually-annotated frames. We use this initial model to estimate the 3D coordinates of the body surface points, and find the correspondences between the triangulated 3D points and those estimated by RatBodyFormer by minimizing their Euclidean distance [26]. These correspondences allow transfer of marker IDs from the manually annotated body surface points to the automatically triangulated body surface points so that RatBodyFormer can be retrained using both the manually-annotated and automatically-annotated frames. We define this automatically-annotated labels as the semi-automatically annotated labels to distinguish them from manually annotated labels. We experimentally show that the use of semi-automatically annotated labels improve estimation accuracy (Sec. 4.2).

### v3

# RatBodyFormer: Rat Body Surface from Keypoints

@@ Abstract
Analyzing rat behavior lies at the heart of many scientific studies. Past methods for automated rodent modeling have focused on 3D pose estimation from keypoints, e.g., face and appendages. The pose, however, does not capture the rich body surface movement encoding the subtle rat behaviors like curling and stretching. The body surface lacks features that can be visually defined, evading these established keypoint-based methods. In this paper, we introduce the first method for reconstructing the rat body surface as a dense set of points by learning to predict it from the sparse keypoints that can be detected with past methods. Our method consists of two key contributions. The first is RatDome, a novel multi-camera system for rat behavior capture, and a large-scale dataset captured with it that consists of pairs of 3D keypoints and 3D body surface points. The second is RatBodyFormer, a novel network to transform detected keypoints to 3D body surface points. RatBodyFormer is agnostic to the exact locations of the 3D body surface points in the training data and is trained with masked-learning. We experimentally validate our framework with a number of real-world experiments. Our results collectively serve as a novel foundation for automated rat behavior analysis.
Figure 1: We introduce a novel multiview camera system (RatDome) to capture multiview videos of rats (top), and a novel Transfomer-based network (RatBodyFormer) that recovers the deforming body surface as a dense set of surface points (bottom: rainbow points) predicted from detectable keypoints (middle: orange points). The body surface reconstructed with RatBodyFormer offers a much richer window into the complex rat behavior.
@@ 1 Introduction
Rodent behavior analysis underpins the scientific process of many areas in biomedical and neuroscientific research. The behavioral outcomes of rodents play a key role in validating hypotheses that lead to scientific discoveries. Careful observations of rodent behaviors, however, incur costs in manpower and time, causing a bottleneck in the scientific process. Human observation is also prone to errors. Subtle behavioral differences may be missed or misjudged [12, 39]. If we can automate this experimental validation process while ensuring its accuracy or even extend it beyond the human level, we may dramatically expedite the scientific discovery process [18, 42, 15, 11].
@@
We focus on rats. Rats are the main animal model for neuroscientific experiments, due to the complexity of their behavioral repertoire and their body size, which is suitable for measurement systems such as electrophysiology [29, 32, 23]. A growing number of recent works, indeed, target computational rat behavior analysis or provide the basis for it. Maghsoudi et al. [41] painted markers on the body and tracked them based on their colors. Matsumoto et al. [44] tracked rat body parts by fitting a physical body model to the point clouds. More recently, Mathis et al. [1] introduced a pre-trained deep network that can be fine-tuned to extract 2D keypoints of interest. This provides a general framework for skeleton estimation of animals, very much in the spirit of human pose estimation [27]. Following the work of Mathis et al., many works have achieved deep learning-based pose estimation of rats in 2D [60, 57] and in 3D [14, 59] from videos or images.
@@
These past methods, however, are pose estimation and only provide us with the means to understand rat behavior through the movements of sparse keypoints that are well-defined visually. With DeepLabCut [1], these keypoints for a rat would typically lie on the face, hands, feet, and tail. Although the movements of these keypoints may suffice for a range of applications, for instance, to recognize when a rat is feeding itself, they only give an extremely sparse sampling of the entire body which tells us much less than what the detailed movements of the body may convey.
@@
The largest missed opportunity lies in the body surface, the fur-coated elastic body that shows subtle twitches, twists, curl-up, stretching, and even hair standing which all eloquently speak to the rich inner conditions of the rat and colorize their behavior with context. Accurate reconstruction of the body surface is difficult with a keypoint-based method, even with large-scale pre-training, as the rat body lacks any visual features whatsoever for such keypoint detectors to latch on to. It is also highly non-rigid and deformable, much more so than the human body. As such simple interpolation of detectable keypoints cannot work.
One may imagine a few approaches for reconstructing this featureless, non-rigid body surface. For chimpanzees, Sanakoyeu et al. [53] successfully leveraged DensePose [20] by transferring the surface points on the chimpanzee body onto the canonical uv-textured surface. This reuse of DensePose, however, is only possible for proximal animal classes of humans as they have the same body structure. For other classes of animals, we are forced to restart from training DensePose itself, which necessitates massive pixel-to-surface annotations. This is not only impractical but infeasible as there is no principled means to define a UV coordinate on the featureless and highly deformable rat body. Instead, Bolaños et al. [8] used a 3D virtual mouse mesh model by CT scanning an anesthetized mouse. MAMMAL [2] tracks the body of a mouse by deforming this same mouse mesh to align with detected keypoints and silhouettes of multi-view images. The body of an awake rat is, however, completely different from an anesthetized one, let alone a rat is very different from a mouse, and cannot be CT-scanned.
@@
How then can we accurately model the 3D body surface deformation of rats? Our goal is to model the non-rigid surface deformation and realize full 3D rat reconstruction including the body surface without interfering with its natural behavior. We achieve this by learning to predict a dense set of body surface points from the detectable facial and appendage keypoints. In other words, our method learns the coordination between the body surface and other body parts namely the head, feet, and tail.
@@
We make two key technical contributions for this. The first is a novel means to collect training data. We need a sizable amount of training data that associates representative keypoints including facial features and appendages with densely sampled points of the body surface. The rat body surface is featureless and deformable, to the point that it is near impossible even for a human to annotate. We overcome this by temporarily attaching trackable points (colorful beads) and passively capturing them in a novel multi-camera system which we refer to as the RatDome. We show that we can automatically recover the keypoints together with the 3D body surface points at each time instance with multiview geometry. We call this first-of-its-kind dataset, the RatDome dataset. The second is a novel network to transform detected 3D keypoints to 3D body surface points. We realize this with a transformer-based model that is agnostic to the exact locations of the 3D body surface points in the training data and introduce masked-learning to train it. We refer to this model as RatBodyFormer.
@@
We experimentally validate our framework with a number of real-world experiments using rats for neuroscientific studies. We first show that RatBodyFormer can estimate the body surface accurately regardless of their poses or shapes. Our model can estimate their body surface with an average L2 error of around 6.5 mm, approximately the same diameter as the beads (5.5 mm) attached to the body surface. We also achieved lower errors of body surface points compared to MAMMAL [2]. We demonstrate the use of the recovered body surface by constructing an animatable rat model using Gaussian Splatting [31], which we refer to as GaussianRat. GaussianRat can be rendered in arbitrary pose driven by the keypoints, which can be leveraged for analysis-by-synthesis in downstream applications. We believe our RatDome dataset and RatBodyFormer offer foundational tools for advancing the automation of rat behavior analysis and together open a new avenue of research towards computer vision for science.
@@ 2 Related Work
Markerless, non-invasive methods have also been proposed. Belongee et al. [3] proposed the particle filter based algorithm to track contours of mice. Chaumont et al. [10] tracked mice by modeling the body parts using geometric primitives and linked them with physical constraints. These methods do not recover the body surface itself.
Recently, deep learning based methods [48, 4, 1, 59, 14, 17, 57, 24, 19, 60] have become popular for detecting keypoints. DeepLabCut [1] detects user-defined keypoints from monocular images based on the network of 2D human pose estimation [27]. DANNCE [14] and FreiPose [59] integrate multiview images to estimate 3D keypoints. LiftPose3D [17] achieves monocular 3D animal pose estimation by leveraging a deep neural network for lifting 2D human poses to 3D [43]. All of these methods can only reliably detect sparse keypoints ( e.g. facial points, paws, midline) with well-defined visual features, or interpolate them on the bone which is rigid. The body surface, however, cannot be localized by these methods because it has no easily identifiable feature of color or shape and as it is highly non-rigid. We reconstruct the body surface passively without any markers by modeling the relationship of detectable keypoint positions and the body surface deformation.
A few recent works have targeted rodent body surface modeling [7, 2, 8]. Bolaños et al. [8] constructed a 3D virtual mouse mesh model from CT data. To generate synthetic training data for pose estimation, they deform the scanned anesthetized body surface by articulating the bones. MAMMAL [2] tracked a mouse body surface and extremities in 3D by fitting the same mouse mesh model [8] so that it aligns with the detected 2D keypoints and the silhouettes in multiview images. The deformation of the anesthetized mouse surface, however, cannot accurately model the surface deformation of an awake rat. We capture the body surface of an awake rat by attaching beads and capturing it from multiple views, and learn to predict those coordinates from detectable keypoints.
@@
One common approach for 3D human reconstruction is to learn a statistical model from pre-acquired 3D scans to better condition the solution space [38, 61, 62, 37]. SMPL [38] is a statistical 3D human linear blended shape model parameterized by the person’s shape and pose. The pose parameters are the joint rotations and the shape parameters are the PCA coefficients of a large collection of aligned body scans. SMAL [61] extends this to quadrupled animals. Most 3D reconstruction methods regress these parameters directly from the image with a neural network [30, 58] or optimize their parameters so that projected keypoints align with their imaged ones [5, 26, 34, 25]. There is no such statistical model of rats, mainly because there is no large-scale collection of 3D scans aligned with a canonical 3D surface. A straightforward approach for building such 3D surface dataset is to capture the 3D surface of the target, i.e., rats, first, and align them w.r.t. a canonical 3D scan [6, 13]. This approach, however, cannot provide reliable alignments because most of the points on the rat body surface cannot be localized by their 3D shape or surface color.
A few works directly regress the 3D body surface from images [52, 53, 20]. AnimalAvatars [52] regresses continuous surface embedding [47] from an image to associate each pixel with a point on the target 3D shape in its canonical pose. This CSE estimation, however, requires large-scale dense annotation between 2D image pixels and a 3D canonical surface ala DensePose [20], which is costly and unreliable for featureless body surfaces, or requires the target animal to be in a proximal class of humans [53]. This annotation process is not only labor-intensive but also limits the accuracy of the 3D surface estimation since the mapping between each pixel and the surface point is determined intuitively by human annotators. We instead leverage beads and multiview geometry to learn mapping from sparse keypoints to dense body surface points, which ensures robust positional consistency. We use beads instead of fluorescent dye [9], since they are easier to distinguish by color.
@@ 3 Method
We reconstruct the 3D body surface of a freely moving rat by learning the mapping from the 3D positions of visually well-defined keypoints, i.e., face and appendages, to the 3D coordinates of a dense set of body surface points. First, we develop a novel multiview camera system which we refer to as the RatDome (Sec. 3.1) and collect a first-of-its-kind dataset that associates densely sampled body surface points with detectable keypoints (RatDome Dataset). Second, we derive a transformer-based network, RatBodyFormer, that takes the keypoint 3D coordinates as input and outputs sampled body surface point 3D coordinates (Sec. 3.2). The model is trained on the RatDome Dataset.
@@ 3.1 RatDome Dataset
Regardless of the pre-training, a network can only learn to detect points that exhibit sufficient visual features. As a result, even after fine-tuning, an off-the-shelf deep keypoint detector [1] can only reliably detect points on the face, appendages, and midline of the rat. Our goal is to learn to extrapolate the body surface from these keypoints. For this, we will need a sufficiently large-scale dataset of paired sets of 3D keypoints and 3D body surface points. This is challenging as the rat body is completely featureless.
We overcome this challenge by simply endowing the rat body with visual features. As shown in Fig. 2, we attach color beads to the body as well as paint markers in the areas where beads cannot be attached. We refer to these two types of body markers, beads and paints, simply as markers. It is important to note that these markers are only used for training data capture and the rats are completely in their natural form when their behavior is observed in actual experiments. These markers are also small enough that they do not alter the rat’s behavior as far as we could tell.
Figure 2: A color-beaded rat (left) and RatDome (right). We attach color (red, black, orange, blue) beads and paint on the rat body surface . RatDome is a novel multiview camera studio for freely moving rats. It is shaped as a 15-faced gyroelongated pentagonal pyramid. With 15 cameras and their multiview geometry, we collect, annotate, and reconstruct paired sets of 3D keypoints and 3D body surface points of rats of different ages in weeks.
We build a novel multi-camera system to passively observe the color-beaded rat and reconstruct the 3D coordinates of individual markers. As shown in Fig. 2, this RatDome is a rat-scale multiview studio. Its shape is a gyroelongated pentagonal pyramid with 15 faces. Each face is an equilateral triangle with sides of 400mm, and accommodates up to three cameras or microphones on it while also serving as a green background of the system. For our data capture, we mount one camera on each face totaling 15 views. RatDome follows the modular design of CMU Panoptic studio [28] and can adapt to new capture devices by replacing each face.
By capturing freely moving rats of different ages (7, 9, 11 week-old) in RatDome, we collect multiview videos and the paired sets of 3D keypoints and 3D markers. We recorded 2 sessions for the 7 week-old, 1 session for the 9 week-old, and 2 sessions for the 11 week-old. Each session was approximately 10 to 15 minutes. We manually annotate the marker IDs for the 7 and 11 week-old rats for about 1100 frames each. We semi-automatically annotate (see Sec. 3.2) the 7, 9, and 11 week-old rats’ marker IDs for about 9000, 1800, and 15000 frames, respectively. The 7, 9, 11 week-old rats have 58, 33, and 66 surface markers, respectively. We refer to this first-of-its-kind large-scale rat body surface dataset as the RatDome Dataset.
@@ 3.2 RatBodyFormer
RatBodyFormer takes $K$ 3D keypoint coordinates as input and outputs $N$ 3D surface point coordinates. The keypoints are chosen as points locatable in images, and we use the following 10 points: nose, right and left eyes, ears, front paws, back paws, and the base of tail.
@@
Figure 3: RatBodyFormer is an encoder-decoder Transformer model that takes the normalized displacements of detected 3D keypoints and outputs the normalized displacements of densely sampled 3D body surface points. The displacements are w.r.t. the reference pose.
@@
To represent different body shapes of different rats, we introduce point-wise scaling and translation parameters for individual rats, $\mathbf{C}_{\mathrm{P}}=\{c_{p_{i}}\}_{1}^{K},\mathbf{T}_{\mathrm{P}}=\{t_{p_{i}}\}_{1}^{K}$ and $\mathbf{C}_{\mathrm{B}}=\{c_{b_{j}}\}_{1}^{N},\mathbf{T}_{\mathrm{B}}=\{t_{b_{j}}\}_{1}^{N}$ for keypoints and body surface points, respectively. These scaling and translation parameters are optimized in the training process. At inference time for a rat whose scaling and translation parameters are unknown, we test-time optimize the parameters while keeping RatBodyFormer frozen.
The keypoints and the body surface points are first rotated around the Z-axis, i.e., the direction of gravity, such that the rat face is oriented to a pre-determined direction. Their 3D coordinates are then normalized to $[-1,1]$: $\mathbf{P}=\{p_{i}\}_{1}^{K}$ and $\mathbf{B}=\{b_{j}\}_{1}^{N}$. The same normalization is also applied to $\tilde{S}$.
As shown in Fig. 3, the decoder outputs are the estimated normalized displacements of the body surface points from their coordinates in the reference pose. The decoder queries are normalized displacements of the body surface points in an initial guess $\hat{\mathbf{B}}$ from their coordinates in the reference pose. The initial guess is calculated analytically. We use ARAP [55] and deform $\tilde{S}$ with the constraint that $\tilde{\mathbf{P}}$ is aligned with $\mathbf{P}/\mathbf{C}_{\mathrm{P}}$. As a result, we obtain the ARAP-deformed $\tilde{\mathbf{B}}$ as $\hat{\mathbf{B}}=\{\hat{b}_{j}\}_{1}^{N}$. For each keypoint and body surface point, we use the displacement vector normalized by the identity-dependent parameters $p_{i}/c_{p_{i}}-(\tilde{p}_{i}+t_{p_{i}})$ and $\hat{b}_{j}-(\tilde{b}_{j}+t_{b_{j}})$ as the inputs to the model, respectively.
The encoder and the decoder first embed each of these displacements in higher dimension $d$ together with positional encoding. The positional encoding for keypoints and body surface points is defined with a set of sinusoidal functions of the projection of their positions $\tilde{p}_{i}$ and $\tilde{b}_{j}$ on $\tilde{S}$ by the Laplace-Beltrami eigenfunctions defined on the canonical body surface $\tilde{S}$ [47, 36]. The outputs of the RatBodyFormer decoder are $N$ normalized displacement vectors $\{\delta_{j}\}_{1}^{N}$. Each $\delta_{j}$ reconstructs the body surface coordinates as $c_{j}(\delta_{j}+\tilde{b}_{j}+t_{b_{j}})$. We set $d=128$ in the experiment.
@@
The primary loss function for training RatBodyFormer is the supervision provided by the annotations in the RatDome Dataset. We directly compare the estimated 3D body surface point coordinates and their ground-truth coordinates with the L2 norm as $\mathcal{L}_{\mathrm{3D}}$. This L2 norm loss $\mathcal{L}_{\mathrm{3D}}$, however, can be defined only for a subset of the $N$ body surface points. Since they are defined as the union of all annotated body surface points on different rat surfaces, a single input keypoint set from a rat has a ground-truth annotation only for the body surface points of the same rat. To ensure that the body surface point positions are consistent, we introduce a position consistency loss $\mathcal{L}_{\mathrm{p}}$ when training on data of multiple rats. Please refer to the appendix for details. In addition to $\mathcal{L}_{\mathrm{3D}}$ and $\mathcal{L}_{\mathrm{p}}$, we employ a silhouette loss $\mathcal{L}_{\mathrm{s}}$ which counts the number of predicted body surface points whose projections fall outside of the 2D rat region in a view [21].
@@
The RatDome Dataset contains many frames capturing color-beaded rats but without manual annotations (Sec. 3.1). We derive a semi-supervised learning method to make full use of this data. The method leverages a small number of frames fully-annotated with ground-truth to infer the labels for the remaining frames.
First, we compute the 3D coordinates of the markers. This process is automated, similar to that of the Panoptic Studio Dataset [28]. We detect the 2D markers from each image with an object detector [16] trained with manual annotations. We then leverage multiview triangulation to disambiguate markers of the same color on the body surface.
Once the 3D marker coordinates are computed for each frame, we assign their marker IDs so that they are consistent with the manually annotated ones by using RatBodyFormer itself. Suppose we have trained RatBodyFormer using only the manually-annotated frames. We use this initial model to estimate the 3D coordinates of the body surface points, and find the correspondences between the triangulated 3D points and those estimated by RatBodyFormer by minimizing their Euclidean distance [35]. These correspondences allow transfer of marker IDs from the manually annotated body surface points to the automatically triangulated body surface points so that RatBodyFormer can be retrained using both the manually-annotated and automatically-annotated frames. We define these automatically-annotated labels as the semi-automatically annotated labels to distinguish them from manually annotated labels. We experimentally show that the use of semi-automatically annotated labels improve estimation accuracy (Sec. 4.2).